OpenStudy (anonymous):

How do I write a linear growth function that passes through two points, and an exponential growth function that will pass through one of the same points. The example I am using goes through points (-5,6) and (-8,31). Can anyone tell me how to set this up? This is algebra 1. Thank you in advance!

OpenStudy (anteater):

To find a linear growth function using two points, you can use the two given points to find the slope and then use the slope and one of your points to create the equation of the line.

OpenStudy (anteater):

Slope (m) is given by "rise"/"run", that is the vertical change divided by the horizontal change. Or (y2 - y1)/(x2 - x1)

OpenStudy (anteater):

So let one of your points be (x1,y1) and the other be (x2,y2). It doesn't matter which.

OpenStudy (anteater):

If we let (-5,6) be (x1,y1) and (-8,31) be (x2,y2), then (y2 - y1)/(x2 - x1) = (31 - 6)/(-8-(-5)) which is 25/-13 or - 25/13

OpenStudy (anonymous):

I did that and got 31-6/-8-(-5) = 25/13 = 1.92, so I am track with that, but it all falls apart after that.

OpenStudy (anteater):

I think we can get it back on the rails :D

OpenStudy (anteater):

Now, you have a couple of choices. You can use either the point-slope formula or the slope intercept form to come up with the equation of your line.

OpenStudy (anteater):

Do you have a preference?

OpenStudy (anonymous):

slope intercept I think Whichever is easier!

OpenStudy (anteater):

I think it may be easier to start with point-slope, then rearrange the equation to put it in slope intercept form.

OpenStudy (anonymous):

I think the slope is 25/-3? -8-(-5)=-8+5=-3?

OpenStudy (anteater):

Yes :) You are correct.

OpenStudy (anteater):

I am addition challenged this morning!

OpenStudy (anonymous):

I am so glad! This is part of a big project and I can't figure out this portion. It is so frustrating!

OpenStudy (anteater):

So we could use point slope: y - y1 = m(x - x1), since we have m and can use either point.

OpenStudy (anonymous):

ok

OpenStudy (anteater):

Just plug them into the formula. Let's use the first point (-5,6)

OpenStudy (anteater):

So we have y - 6 = (-25/3)(x - (-5))

OpenStudy (anonymous):

I am with you

OpenStudy (anteater):

And then we distribute the -25/3 on the right side: y - 6 = (-25/3)x -125/3

OpenStudy (anteater):

Making sense so far? They gave you ugly numbers to work with!

OpenStudy (anonymous):

I know, my other set of numbers is even worse! I have to do this twice, so I put the easier #'s here, and then I can figure out the other set once I have an example.

OpenStudy (anteater):

Makes sense to me!

OpenStudy (anonymous):

I had to make up a parabola and certain points, so that is where the #s came from.

OpenStudy (anteater):

Aaah

OpenStudy (anteater):

So now you do have the equation of a line in point slope form. If you want it in slope intercept (y = mx + b) form, you just need to isolate y. So then just add 6 to both sides.

OpenStudy (anonymous):

Ok, any thoughts on an exponential. I watched a few videos, but can't figure it out. The lesson gives equations, but doesn't explain how to write them.

OpenStudy (anteater):

y = (-25/3)x - 125/3 + 6 and then combine the -125/3 + 6

OpenStudy (anteater):

Yes, for exponential you also use both points.

OpenStudy (anteater):

But here you start with the general equation: y = ab^x

OpenStudy (anteater):

And the trick is to figure out what a and b are.

OpenStudy (anteater):

So, you start by creating two equations using the points you were given. So, the first one would look like: 6 = ab^(-5)

OpenStudy (anteater):

And so the second one would be ...?

OpenStudy (anonymous):

The second one is 31=ab^-8

OpenStudy (anteater):

Of course. :) And then the easiest thing to do is usually to find b first by eliminating a.

OpenStudy (anonymous):

That is were I was having trouble.

OpenStudy (anteater):

So, one way you could do that is to divide the second equation by the first. In that way the a's will cancel out.

OpenStudy (anteater):

So, then 31/6 = (ab^-8)/(ab^-5)

OpenStudy (anteater):

What would that simplify to then?

OpenStudy (anonymous):

The a's cancel and you have b^-8 over b^-5 on the right side? Left is 31/6.

OpenStudy (anteater):

Yes. And then b^-8 divided by b^-5 simplifies to ...

OpenStudy (anonymous):

I forgot the rule, trying to look back, the two negatives are confusing me.

OpenStudy (anteater):

That's ok. I won't make you look it up! When you divide, you subtract the exponents, so -8-(-5)

OpenStudy (anonymous):

-3 again!

OpenStudy (anteater):

Yep!

OpenStudy (anteater):

So we have 31/6 = b^-3

OpenStudy (anteater):

Or 31/6 = 1/b^3

OpenStudy (anteater):

Because of the reciprocal property of negative exponents

OpenStudy (anonymous):

right, I forgot about that!

OpenStudy (anteater):

So then, b^3 = 6/31

OpenStudy (anteater):

And then b will be the lovely cube root of 6/31 !!!

OpenStudy (anteater):

Which is just heinous, if you ask me!

OpenStudy (anonymous):

ugh, I couldn't have picked worse numbers

OpenStudy (anteater):

Oh, I bet you could have, actually :D

OpenStudy (anonymous):

That is part of why it was so hard to figure out, the samples all worked out to squares. So how do I work that into an exponential equation.

OpenStudy (anteater):

You just leave it as it is written. Put it in parentheses, then raise it to x.

OpenStudy (anteater):

One way you could write it is b = (6/31)^(1/3)

OpenStudy (anonymous):

It has to make an exponential line with geogebra, so I am not sure if that format works.

OpenStudy (anteater):

So then when you raise it the the x power, you would have (6/31)^(1/3)*x

OpenStudy (anteater):

You know, my son used geogebra before

OpenStudy (anteater):

So I think we can get it formatted properly at the end, after we figure out what a is.

OpenStudy (anonymous):

I think it needs to be a function, like f(x)=

OpenStudy (anteater):

Yes, I believe you are correct. So we just need to find our a. Then our function will be f(x) = a*(6/31)^(x/3)

OpenStudy (anonymous):

Ugh, it says illegal multiplication. Let me try it a different way.

OpenStudy (anteater):

What did you get for a?

OpenStudy (anonymous):

I put it in with a? I don't know?

OpenStudy (anteater):

We need to get a first. I think that is why it said "illegal multiplication", since we need to put a number in for a.

OpenStudy (anteater):

So, now that we have b, we can use y = ab^x and one of the points to go back and find a.

OpenStudy (anteater):

It won't be a pretty number. ;)

OpenStudy (anonymous):

Oh, this is supposed to be a fairly simple project, I have spent days on it!

OpenStudy (anteater):

We can use (-5,6) again, if you want?

OpenStudy (anonymous):

Sure

OpenStudy (anteater):

I am sorry to hear that! I think you are right, the choice of points makes a big difference to how easy the problem is.

OpenStudy (anteater):

Generally speaking, if you can choose points where x = 0 and where x = 1

OpenStudy (anteater):

that will simplify your calculations

OpenStudy (anteater):

Do you have an image of the original question?

OpenStudy (anteater):

What you are doing should "work", it's just not the easiest way on account of the numbers.

OpenStudy (anonymous):

I should all work, it just isn't easy, lol.

OpenStudy (anonymous):

It not I

OpenStudy (anonymous):

yes, hold on

OpenStudy (anteater):

LOL

OpenStudy (anteater):

I should work too, but this is me procrastinating! ;)

OpenStudy (anonymous):

West Mathington’s most urgent need is a parabolic freeway. Create your own upward opening quadratic function, f(x), which has two real zeros. Prove that f(x) has two real zeros. If you are doing this project collaboratively, then your partner will need to create a unique second freeway function, g(x), and prove that it also has two real zeros. Two on-ramps are needed to be placed on the parabolic freeway. Decide where on the parabola of f(x) you are placing the on-ramp locations. Write those ordered pairs down. If you are working collaboratively, identify four on-ramps, two on each function. West Mathington wants to connect these on-ramps with some surface roads. Create a linear growth function, h(x), that passes through both on-ramp points. Create an exponential growth function, j(x), that passes through at least one of the on-ramp points. Show all of the work you did to create both functions. Collaborative pairs need to connect linear function h(x) through the on ramp points on f(x), and connect exponential function j(x) through any on ramp point on g(x). At least three of the four on ramp points will be used. What important relationship does the x-coordinates of the on-ramp location points have with the system of equations formed by the two roads’ functions that are being connected? Provide justification and support for your explanation. The city planner needs to identify the most northern road. Prove which road will eventually go the furthest to the north (positive y-direction). Create tables for your functions using an appropriate domain of five integers. Using the tables and graph, explain to the city planner which road will be the furthest north as the x values continue to get larger (the road continues to go east). Provide reasoning why. Collaborative groups should prove whether f(x) or h(x) is the most northern between the two, and whether g(x) or j(x) is the most northern between those two. Include your graph that shows the functions that model each of the roads and the on-ramps.

OpenStudy (anteater):

If you could upload an image of the original problem, we could probably find two different points and come up with linear and exponential equations quickly. Or, we can finish finding a for the points we were using and then go ahead and submit the resulting function.

OpenStudy (anonymous):

I have to do it as a collaboration, so I did 2 parabolas, then picked 4 points, now on the part with linear growth for both on ramps and exponential growth.

OpenStudy (anonymous):

The posts are out of order on my screen, do you see it?

OpenStudy (anonymous):

I didn't want to redo the whole thing, but I am open to suggestions at this point. I don't know why I am having such a hard time with this.

OpenStudy (anteater):

Yes, it is very confusing!

OpenStudy (anteater):

Oh, it is way up at the top! Just a moment, please, and I'll read it. :)

OpenStudy (anteater):

Wow.

OpenStudy (anteater):

Can I ask you a question?

OpenStudy (anteater):

What class are you taking and in pursuit of what major? What book is this from?

OpenStudy (anteater):

I can tell you why you're having a hard time with it. It's a pretty wild problem!

OpenStudy (anteater):

Are we allowed to communicate outside of OpenStudy? I don't know what the "rules" are regarding privacy of users, but I think we could do this more easily by email!

OpenStudy (anonymous):

Sorry my computer froze. IDK the rules, but If you want my email I am not sure how to do that without giving it to everyone

OpenStudy (anonymous):

This is an FLVS honors algebra class. So yeah, 8th-9th grade!

OpenStudy (anonymous):

Sorry, FLVS is Florida Virtual School.

OpenStudy (anonymous):

I am trying to help my daughter with this. . .